Calculus I, Fall 2014 (http://www.math.nthu.edu.tw/~wangwc/)

Brief answer to selected problems in Homework 13

1. Section 6.4: problem 22: S = ∫_{0}^{√\overline{2}} 2πx\sqrt{1 + x^{2}(x^{2} + 2)}dx = 4π